OpenStudy (anonymous):

x-2(x+10)=12 please explain

TheSmartOne (thesmartone):

Distribution property. a(b-c)=ab-ac

TheSmartOne (thesmartone):

-2(x+10)=?

OpenStudy (anonymous):

huh?

TheSmartOne (thesmartone):

So -2(x+10)=-2(x)+(-2)(10)=??

OpenStudy (arabpride):

Distribute: x - 2 ( x + 10 ) = 12 x- (2*x) + (2*10) = 12 x - 2x + 20 = 12 Then you combine like terms~ Do you know how to do that? :D

OpenStudy (anonymous):

@arabpride not really im kind of slow with it

TheSmartOne (thesmartone):

What grade are you in? If I may ask?

OpenStudy (anonymous):

8th why

OpenStudy (arabpride):

The good times *-*

TheSmartOne (thesmartone):

:)

OpenStudy (arabpride):

Sorry - i was afk >.> okay now ill help u finish :D

OpenStudy (anonymous):

@arabpride it okay i understand

TheSmartOne (thesmartone):

x - 2x - 20 = 12 Subtract like terms You get -1x-20=12 Add 20 on both sides -1x - 20 = 12 +20 +20 -------------------- -1x =32 x=-32

OpenStudy (arabpride):

Combine like terms: just basically means to put together common variables and whole numbers etc.~ x - 2x + 20 = 12 -20 -20 x - 2x = -8 -x = -8 we need a POSITIVE answer so we multiply BOTH sides be -1 (-1)*-x = (-1)*-8 x = 8 :D

TheSmartOne (thesmartone):

lol The conflicting answers.

OpenStudy (arabpride):

HOLY MOLY 2 different answers x'DD

TheSmartOne (thesmartone):

Les plug it in :)

OpenStudy (arabpride):

WAIT!!

OpenStudy (arabpride):

It's -3x!!! <phacewack> x-x

OpenStudy (arabpride):

Okay lemme correct myself >.>

TheSmartOne (thesmartone):

x-2(x+10)=12 -32- 2(-32+10)=12 -32-2(-22)=12 -32+44=12 12=12

TheSmartOne (thesmartone):

Yay I was correct :)

TheSmartOne (thesmartone):

Look at us. We are mathlete' but are stumbling on this problem.. lol

OpenStudy (anonymous):

wait im confused who has the right answer?? :/

OpenStudy (arabpride):

Wait where did 32 come from? o-o

TheSmartOne (thesmartone):

x - 2x - 20 = 12 Subtract like terms You get -1x-20=12 Add 20 on both sides -1x - 20 = 12 +20 +20 -------------------- -1x =32 x=-32

TheSmartOne (thesmartone):

I just googled it and look what I found http://www.geteasysolution.com/x-2(x+10)=12

TheSmartOne (thesmartone):

SO I was correct.

TheSmartOne (thesmartone):

So*

TheSmartOne (thesmartone):

Note to self: Always Google questions before attempting them :)

OpenStudy (arabpride):

No no - u cant ADD the 20 because it's A POSITVE - u have to SUBTRACT it~

OpenStudy (arabpride):

unless -2 distributed... /.\

TheSmartOne (thesmartone):

Yes but \(\Huge x\color{red}{-}2(x+10)=12 \)

OpenStudy (anonymous):

this one is hard

OpenStudy (arabpride):

Yah LOLZ i forgot that 2 was negative - okay yahh i was wrong -sorry - threw u off cx

TheSmartOne (thesmartone):

You forgot to distribute the - sign..

TheSmartOne (thesmartone):

SO yeah :)

TheSmartOne (thesmartone):

So*

OpenStudy (arabpride):

It really isnt - ive just been getting into higher math and my short term memory :P

OpenStudy (arabpride):

smh~

OpenStudy (anonymous):

lol thank u both for ur help so the answer is 32 right

TheSmartOne (thesmartone):

\(\Huge -32\)

OpenStudy (arabpride):

wrote it down just now ~ x = -12 right?

TheSmartOne (thesmartone):

Can you read that.. lol

OpenStudy (arabpride):

OMG AGAIN!??!?! LOLZ!!!

OpenStudy (arabpride):

WAIT NVM!~

TheSmartOne (thesmartone):

Haha Trust me I am a genius.. 14 year old in 11th grade.

OpenStudy (arabpride):

OMg smh~ forget it :'D this is too funneh - i cant even rn~

TheSmartOne (thesmartone):

:)

OpenStudy (anonymous):

lol @TheSmartOne u r smart

OpenStudy (arabpride):

I coulda grasuated by 15 >.> but :/ long story

TheSmartOne (thesmartone):

I had this before you came @arabpride but all is well xD :D :)

OpenStudy (arabpride):

graduated**

OpenStudy (anonymous):

u r too @arabpride

TheSmartOne (thesmartone):

Thanks @lelelove

OpenStudy (arabpride):

Yea - sorry - LOLZ yahh math is my least favorite subject tbh~ i prefer history but thanks *^-^

TheSmartOne (thesmartone):

I will graduate High School when I am 16 with an Associates Degree xD xD :)

OpenStudy (arabpride):

Lucky You :D im gunna graduate and sit at home and do nuthing :P

TheSmartOne (thesmartone):

I don't understand how you use all those emoticons/symbols and so many of them at once lol

OpenStudy (arabpride):

So there's no point in trying to graduate early for meh x-x

OpenStudy (arabpride):

LOLZ i just doo x'DD do i over do them? o-o

OpenStudy (anonymous):

@TheSmartOne WOW thats good i feel slow now lol @arabpride lol same here

TheSmartOne (thesmartone):

I gotta graduate and get a good career.. Gotta do good in the SAT, get into a good college and live a happy life ever after.

TheSmartOne (thesmartone):

BTW I like both of your profile pics.

TheSmartOne (thesmartone):

Your welcome xD

OpenStudy (arabpride):

I had it all planned out too tbh... but i realized i got my hopes up :/ A Career isnt exactly an option for meh ._.

OpenStudy (anonymous):

which one of r u talking to? and lol @TheSmartOne

TheSmartOne (thesmartone):

I said both :)

TheSmartOne (thesmartone):

So that means @arabpride and @lelelove

OpenStudy (arabpride):

Lolz thanks *^-^

TheSmartOne (thesmartone):

Join in @Nnesha

OpenStudy (arabpride):

But srsly do u think i overdo the emoticons? o-o

TheSmartOne (thesmartone):

I just realized this is still open.. lol

TheSmartOne (thesmartone):

Maybe @arabpride :) But its cool..

OpenStudy (arabpride):

<.< thanks fer letting meh know cx omg cant even help myself anymore x'DD

TheSmartOne (thesmartone):

Did you see Compassionate's profile? He is trying to make people flirt with him so he can report them lol

TheSmartOne (thesmartone):

http://openstudy.com/users/compassionate

TheSmartOne (thesmartone):

:)

OpenStudy (arabpride):

OMG LOLZ!!!!! Man i shuda reported all the d00ds who tried to flirt with meh x'DD but i report no one tbh~

TheSmartOne (thesmartone):

xD its hilarious change your pic to something like that and all these users who have pics of them showing their 6-packs start messaging you.. lol

OpenStudy (anonymous):

LOL

TheSmartOne (thesmartone):

Man I need to work hard to get to 99 *sigh*

OpenStudy (arabpride):

ive changed it SOO much but i thnk its some1 trolling meh cx at one point it was SO bad i put a pic of my brother up - smhu can guess: SO.MANY.GIRLZ!!!

OpenStudy (anonymous):

it shouldnt be hard for u

TheSmartOne (thesmartone):

lol @arabpride

OpenStudy (anonymous):

@arabpride lol

OpenStudy (arabpride):

it probably isnt some1 was able to do it in a month x'D

TheSmartOne (thesmartone):

It takes a while to get to 65 and then to 90

TheSmartOne (thesmartone):

I know but I don't go to an online school so I am not infront o a computer all ay long.

OpenStudy (arabpride):

Alright being taggged im off this post - later peeps cx

TheSmartOne (thesmartone):

Oh yeah @Nnesha Congrats on passing 75 SS and join the puke color xD

TheSmartOne (thesmartone):

See ya @Nnesha We need to get to 99 lol

TheSmartOne (thesmartone):

LEts stop spamming this post.

TheSmartOne (thesmartone):

Bye @arabpride @Nnesha @lelelove and not but least @TheSmartOne

OpenStudy (anonymous):

Bye @TheSmartOne

TheSmartOne (thesmartone):

Oh well.